# Powering In-Database Dynamic Model Slicing for Structured Data Analytics

### v1

@@ Abstract.
Relational database management systems (RDBMS) are widely used for the storage and retrieval of structured data. To derive insights beyond statistical aggregation, we typically have to extract specific subdatasets from the database using conventional database operations, and then apply deep neural networks (DNN) training and inference on these respective subdatasets in a separate machine learning system. The process can be prohibitively expensive, especially when there are a combinatorial number of subdatasets extracted for different analytical purposes. This calls for efficient in-database support of advanced analytical methods.
In this paper, we introduce LEADS, a novel SQL-aware dynamic model slicing technique to customize models for subdatasets specified by SQL queries. LEADS improves the predictive modeling of structured data via the mixture of experts (MoE) technique and maintains inference efficiency by a SQL-aware gating network. At the core of LEADS is the construction of a general model with multiple expert sub-models via MoE trained over the entire database. This SQL-aware MoE technique scales up the modeling capacity, enhances effectiveness, and preserves efficiency by activating only necessary experts via the gating network during inference. Additionally, we introduce two regularization terms during the training process of LEADS to strike a balance between effectiveness and efficiency. We also design and build an in-database inference system, called INDICES, to support end-to-end advanced structured data analytics by non-intrusively incorporating LEADS onto PostgreSQL. Our extensive experiments on real-world datasets demonstrate that LEADS  consistently outperforms baseline models, and INDICES delivers effective in-database analytics with a considerable reduction in inference latency compared to traditional solutions.
@@ 1. Introduction
Relational Database Management Systems (RDBMS) are extensively employed as the primary storage solution for structured data across various applications (Khamis et al., 2020; Nikolic et al., 2020; Li et al., 2019; Rendle, 2013). They serve as a fundamental infrastructure for various domains and are critical to the operation of numerous businesses (Jia et al., 2023; Zheng et al., 2020; He and Chua, 2017). In the contemporary business landscape, structured data analytics via databases has become an indispensable component for driving business growth and success (Jia et al., 2023; Zheng et al., 2020; He and Chua, 2017; Ma et al., 2021; Park et al., 2022). Traditional structured data analytics approaches rely on database-driven filtering or aggregation operations to derive insights. However, these insights only offer a limited statistical view, which often fails to capture the complexity and intricacies of the underlying patterns (Gray et al., 1997; Rajakumari and Nalini, 2014). Fortunately, recent advancements in Deep Neural Networks (DNNs) open up new horizons for advanced analytics beyond simple statistical aggregation (Gardner and Dorling, 1998; Lian et al., 2018; Cheng et al., 2020; Cai et al., 2021).
@@
At its core, exploiting DNNs for advanced structured data analytics comprises two main phases: training and inference (Gardner and Dorling, 1998). The former primarily involves the construction of a DNN model and the training of this model on targeted data, while the latter utilizes the trained model to make predictions on new data. Notably, to deliver advanced DNN-driven analytics for informed decision-making, effectiveness and efficiency are the two most important metrics to optimize for (Davis and Goadrich, 2006; Kwon et al., 2019; Cai et al., 2019; Wasay et al., 2021). Specifically, effectiveness focuses on the inference phase, measuring the extent to which the predictions delivered by the model are accurate. Meanwhile, efficiency evaluates the requirements of the model in terms of response time and computational resources in both phases (Kwon et al., 2019).
@@
In real-world scenarios, analysts are often more interested in performing analytics on specific subsets of data. For instance, they may assess trends among patients diagnosed with a particular disease, or, study behaviors of consumers of a certain age group. Consider the scenario illustrated in Figure 1, where an analyst aims to evaluate the influence of education and city location on the incomes of different subdatasets, i.e., tuples grouped by gender and age. Naturally, the analyst seeks to build a predictive model that is effective, delivering accurate predictions for these subsets of tuples, and meanwhile, executes predictions efficiently with minimal response time and computational resources. However, there are two main challenges in achieving this objective.
Figure 1. An illustration of in-database analytics on income via SQL-aware dynamic model slicing.
@@
First, achieving efficient training for effective predictive modeling across analyst-specified subdatasets is challenging. Conventionally, a single general model is trained to support inference across all data tuples (He and Chua, 2017; Khamis et al., 2020; Gardner and Dorling, 1998). This approach is efficient, which requires training only one model. However, such a model, optimized to capture the common patterns and general behaviors of the whole dataset, is likely not as effective in providing accurate predictions as a dedicated model trained on a specific subdataset of interest. Taking for example the scenario in Figure 1, a model trained explicitly for the group of the gender male and age 24 would probably identify finer-grained patterns and behaviors pertinent to this subdataset, given sufficient training tuples, this dedicated model could outperform the general model significantly. Nonetheless, training a separate model for each subdataset is computationally prohibitive due to the combinatorial nature of potential subdatasets.
Second, efficiently integrating the inference phase into an RDBMS while ensuring effectiveness is also challenging from a system perspective. One major obstacle is how to reconcile the practices of managing structured data within an RDBMS and the execution of inference on a separate ML system. Many existing solutions support the inference process with two separate systems (Wang et al., 2018; Crankshaw et al., 2017), which requires transferring the inference data from an RDBMS, typically a subset of tuples, to another inference system. Such a process is time-consuming, susceptible to errors, and might also violate privacy and security requirements (Xu et al., 2022). Recently, several preliminary attempts have been made to integrate the inference process directly into RDBMS via User-Defined Functions (UDFs) (Abo Khamis et al., 2018; Hellerstein et al., 2012; Feng et al., 2012; Zhang et al., 2021; Yang et al., 2022; Xu et al., 2022), which improves user experience by enabling in-database inference through SQL statements. Specifically, running Python in UDFs can tap into its rich machine learning libraries (Foufoulas et al., 2022), while it misses the opportunity to leverage the more efficient data retrieval APIs offered by RDBMS, e.g., server programming interface (SPI) A proposed enhancement involves utilizing multiple programming languages within UDFs, aiming to harness both data retrieval APIs and advanced ML libraries. However, this approach introduces extra overhead and affects inference efficiency, especially when conversions and copying of inference data between different language execution environments become necessary (Foufoulas et al., 2022). Therefore, achieving efficient and seamless integration of the inference process into RDBMS is an imperative problem to address.
@@
To address the above challenges, we build an efficient and effective IN-Database InferenCE System (INDICES). The system is designed to produce effective predictions across subsets of data dynamically specified and retrieved by SQL queries. To this end, we propose a novel SQL-awarE dynAmic moDel Slicing (LEADS) technique, which enhances the effectiveness of the base model via the mixture of experts (MoE) technique, and maintains the inference efficiency using a SQL-aware gating network for dynamic model customization for subdatasets specified by SQL queries. Specifically, in LEADS, we propose to enhance the modeling capacity of the base model by constructing a general model that is composed of multiple replicas of this base model. These replicas, termed as experts, are trained to specialize in different problem subspaces for more effective predictive modeling. To enhance effectiveness via MoE without incurring reduced inference efficiency, we further introduce a SQL-aware gating network that dynamically generates sparse gating weights based on filter conditions in the SQL query to slice a subset of only necessary experts from the general model. Such a sliced model is optimized for the corresponding SQL query during training, and is dedicated to the specified subdataset for enhancing inference effectiveness while maintaining efficiency.
To support end-to-end structured data analytics, our system INDICES seamlessly incorporates LEADS into PostgreSQL, an open-source RDBMS widely used in both industry and academia. For ease of use and inference efficiency, we divide the proposed in-database inference process into four separate stages and propose three optimization techniques to minimize the overhead of each stage: efficient execution allocation, memory sharing, and state caching. Given that all the stages of the inference process are supported within a single UDF, analysts can now conveniently invoke inference queries using a single SQL statement. This approach obviates the need to transfer and manage data in separate systems and reduce data copying overhead between executions of different programming languages. Additionally, while the current system is supported by PostgreSQL, INDICES can be readily integrated into other RDBMSs, e.g., MySQL.
We summarize the main contributions as follows.
@@
  • We formulate the SQL-aware structured data analytics problem, which requires efficient and effective predictive modeling on subdatasets specified by corresponding SQL queries. To the best of our knowledge, this is the first work that develops techniques and a system to address the problem.
@@
  • We propose a novel SQL-aware dynamic model slicing technique LEADS, which scales up the modeling capacity of the base model via MoE and devises a SQL-aware gating network for efficient and effective dynamic model customization for SQL-specified subdataset.
  • We design and build an end-to-end in-database inference system INDICES for advanced structured data analytics, which non-intrusively incorporate LEADS onto PostgreSQL with three optimization techniques for further improving the inference efficiency.
@@
  • We conduct extensive experiments on four real-world datasets. The results confirm the effectiveness of LEADS, with up to $3.95\%$ improvement in accuracy for given workloads of datasets compared with the baseline models, while INDICES achieves up to $2.06$x speedup in terms of inference efficiency compared with the traditional solution.
@@
In the remainder of this paper, we introduce preliminaries in Section 2. We formulate our problem in Section 3. We present LEADS with detailed descriptions of its modules and optimization schemes in Section 4. We discuss the integration of LEADS and INDICES with PostgreSQL in Section 5. Experimental results are presented in Section 6. We review related work in Section 7 and conclude the paper in Section 8.
@@ 2. Preliminaries
In this section, we present two key techniques central to our system, namely Mixture of Experts (MoE) for scaling up the model capacity while maintaining its inference efficiency via conditional computation (Shazeer et al., 2017), and sparse softmax (Martins and Astudillo, 2016) for the informed selection of active experts for enhancing efficiency. Scalars, vectors and matrices are denoted by $x$, $\mathbf{x}$ and $\mathbf{X}$ respectively.
@@
Mixture of Experts (MoE)  (Jordan and Jacobs, 1994; Waterhouse et al., 1995; Shazeer et al., 2017) is a general ensemble learning and conditional computation technique to scale up the modeling capacity without incurring much computational overhead. In DNN-based MoE, a series of expert DNNs are adopted to divide problem space into different regions, where each expert specializes in handling a certain sub-region. MoE is particularly useful when the data exhibits complex patterns or variations (Riquelme et al., 2021; Fedus et al., 2022; Zhou et al., 2022) due to its capability of enhancing model capacity. There are two main components in an MoE layer: expert models and a gating network. For simplicity of construction, expert models can be composed of homogeneous models that share the same model architecture. During training, expert models are trained to specialize in different problem subspaces. The gating network is also trained to produce a set of gating weights dynamically, which determines the importance assigned to corresponding experts.
@@
Denoting the gating weights and outputs of experts as $\mathbf{w}=[w_{1},w_{2},...,w_{K}]$ and $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2}...,\mathbf{h}_{K}]$ respectively, where $K$ is the number of experts, and $\mathbf{h}_{i}$ is the output of the $i$-th expert, the output of the MoE for the current input is then a weighted average of these experts: $\hat{\mathbf{y}}=\sum_{i=1}^{K}w_{i}\mathbf{h}_{i}$. During training, the MoE model optimizes the gating network and experts simultaneously. The gating network learns to assign appropriate weight to experts, while the experts learn to make accurate predictions within their respective regions of expertise.
@@
MoE has found extensive application in various domains, notably in the large language model GPT-4 (OpenAI, 2023) for texts and the large vision-language model MoE-LLaVA (Lin et al., 2024) for images, which combines the benefits of large model capacity with efficient computation, by only engaging a fraction of the model parameters for each input. In LEADS, we focus on the applicability of the MoE technique to structured data analytics, intending to harness its scalable modeling capacity for enhanced predictive accuracy and efficiency.
@@
Sparse Softmax. Softmax transformation is a crucial function in the gating network, which maps an input vector $\mathbf{z}$ into a probability distribution $\mathbf{p}$ whose probabilities correspond proportionally to the exponential of its input values, i.e., $\mathrm{softmax(z_{j})}=\frac{\mathrm{exp(z_{j})}}{\mathrm{\sum_{i}exp(z_{i})}}$. The output of softmax can thus be subsequently used as the network output denoting the class probabilities or weights indicating the importance of corresponding inputs. Specifically, denoting the $d$-dimension probability as $\rm\Delta^{d}:=\{\mathbf{p}\in\mathbb{R}^{d}:\mathbf{p}\geq 0.\text{ }||\mathbf{p}||_{1}=1\}$, softmax can be interpreted in the variational form with entropy:
(1) $\rm softmax(\mathbf{z})=\mathop{argmax}\limits_{\mathbf{p}\in\Delta^{d}}{\mathbf{p}^{T}\mathbf{z}+H^{S}(\mathbf{p})}$
where $\mathrm{H^{S}(\mathbf{p})}=-\sum_{j}p_{j}\log p_{j}$ is the Shannon entropy. The softmax function is extensively used in DNNs, largely due to its differentiable and convex properties. However, softmax always assigns dense probabilities to inputs, which is less interpretable and effective (Goodfellow et al., 2016; Cai et al., 2021), as compared with sparse credit assignment. To overcome this limitation, sparse softmax is proposed to produce sparse distributions, by assigning zero probability to certain outputs. Particularly, $\alpha$-entmax (Peters et al., 2019) generalizes both dense and sparse softmax with Tsallis $\alpha$-entropies $\rm H_{\alpha}^{T}(\mathbf{p})$ (Tsallis, 1988):
@@
(2) $\rm\alpha\text{-}entmax(\mathbf{z})=\mathop{argmax}\limits_{\mathbf{p}\in\Delta^{d}}{\mathbf{p}^{T}\mathbf{z}+H^{S}_{\alpha}(\mathbf{p})}$
where $\mathrm{H^{S}(\mathbf{p})}=-\sum_{j}(p_{j}=p^{\alpha}_{j})$ if $\rm\alpha\neq 1$, else $\rm H_{1}^{T}(\mathbf{p})=H^{S}(\mathbf{p})$. With a larger $\alpha$, $\alpha$-entmax tends to produce a sparser probability distribution. Another appealing property is that the hyper-parameter $\alpha$, which controls the shape and sparsity of the mapping, can be learned adaptively to the predictive task in the training stage. Let $\mathbf{p}^{*}=\alpha\text{-entmax({z})}$ denote the distribution $\widetilde{p_{i}}=(p^{*}_{i})^{2-\alpha}/\sum_{j}(p^{*}_{j})^{2-\alpha}$ and the Shannon entropy $h_{i}=-(p^{*}_{i})log(p^{*}_{i})$. The gradient of $\alpha$ is derived as: $\frac{\partial\ {\alpha\text{-entmax({z})}}}{\partial\alpha}=\frac{(p^{*}_{i})-\widetilde{p_{i}}}{{(\alpha-1)}^{2}}+\frac{h_{i}-\widetilde{p_{i}}\sum_{j}h_{j}}{\alpha-1},\ \alpha>1$ , which can be optimized end-to-end together with the parameters of the predictive model (Peters et al., 2019). We aim to adopt a learnable sparse softmax in LEADS to improve model training and further enhance the predictive efficiency.
3. Problem Formulation
@@
Technically, structured data can be viewed as one logical table T, which comprises $N$ rows and $M$ attributes within RDBMS. Each row, represented as a tuple $\textbf{x}=(x_{1},x_{2},\cdots,x_{M})$, serves as a feature vector in predictive modeling, with $x_{i}$ denoting the value of the $i$-th attribute. In structured data analytics, data analysts typically focus on specific subsets of data characterized by shared attributes. For example, analysts may assess the readmission rates among the patients diagnosed with a certain disease, or predict the e-commerce click-through rate (CTR) within a particular age group. Typically, for complex analytical queries that involve prediction, WHERE statement in a SQL query is executed first to select relevant tuples, to which DNNs are applied subsequently for prediction. In this paper, we refer to this process as SQL-aware predictive modeling.
@@
Given a SQL query, denoted by $q$, there are two main steps in SQL-aware predictive modeling: data selection and model prediction. Utilizing relational algebra, a generalized SQL query selection $q$ is expressed as $\sigma_{\varphi}(\textbf{T})$, where $\sigma$ is the unary operator for selection and $\varphi$ is the propositional formula in $q$. Typically, $\varphi$ consists of multiple predicates connected by logical operators. The selection $\sigma_{\varphi}(\textbf{T})$ retrieves all tuples in table T that satisfies $\varphi$, formally defined as $\sigma_{\varphi}(\textbf{T})=\{\textbf{x}:\textbf{x}\in\textbf{T},\varphi(\textbf{x})\}$. For simplicity, the subdataset retrieved by the SQL query $q$ is denoted as $\textbf{T}_{\varphi}=\{\textbf{x}_{1},\textbf{x}_{2},\cdots,\textbf{x}_{n}\}$, where $n$ is the number of tuples. Each tuple $\textbf{x}_{i}\in\mathbb{R}^{M}$ in $\textbf{T}_{\varphi}$ comprises $M$ attributes, and $\textbf{x}_{i}$ can be represented as a vector of categorical and/or numerical features, i.e., $\textbf{x}_{i}=[x_{i,1},x_{i,2},\cdots,x_{i,M}]$. DNNs are then applied to perform prediction on these selected tuples, e.g., to predict the labels $\textbf{y}=\{y_{1},y_{2},\cdots,y_{n}\}$, aiming to derive meaningful insights, such as patients readmission rates in healthcare analytics or CTR in e-commerce. Technically, SQL-aware predictive modeling refers to making predictions on a selected subset of tuples retrieved from a logical table T based on a SQL query $q$ with a propositional formula $\varphi$.

### v2

@@ Abstract.
Relational database management systems (RDBMS) are widely used for the storage of structured data. To derive insights beyond statistical aggregation, we typically have to extract specific subdatasets from the database using conventional database operations, and then apply deep neural networks (DNN) training and inference on these subdatasets in a separate analytics system. The process can be prohibitively expensive, especially when there are various subdatasets extracted for different analytical purposes. This calls for efficient in-database support of advanced analytical methods.
In this paper, we introduce LEADS, a novel SQL-aware dynamic model slicing technique to customize models for specified SQL queries. LEADS improves the predictive modeling of structured data via the mixture of experts (MoE) and maintains efficiency by a SQL-aware gating network. At the core of LEADS is the construction of a general model with multiple expert sub-models trained over the database. The MoE scales up the modeling capacity, enhances effectiveness, and preserves efficiency by activating necessary experts via the SQL-aware gating network during inference. To support in-database analytics, we build an inference extension that integrates LEADS onto PostgreSQL. Our extensive experiments on real-world datasets demonstrate that LEADS  consistently outperforms the baseline models, and the in-database inference extension delivers a considerable reduction in inference latency compared to traditional solutions.
@@ 1. Introduction
Relational Database Management Systems (RDBMS) are extensively employed as the primary storage solution for structured data across various applications (Khamis et al., 2020; Nikolic et al., 2020; Li et al., 2019; Rendle, 2013). They serve as a fundamental infrastructure for various domains and are critical to the operation of numerous businesses (Jia et al., 2023; Zheng et al., 2020; He and Chua, 2017). In the contemporary business landscape, structured data analytics via databases has become an indispensable component for driving business growth and success (Jia et al., 2023; Zheng et al., 2020; He and Chua, 2017; Park et al., 2022). Traditional structured data analytics approaches rely on database-driven filtering or aggregation operations to derive insights. However, these insights only offer a limited statistical view, which often fails to capture the complexity and intricacies of the underlying patterns (Gray et al., 1997; Rajakumari and Nalini, 2014). Fortunately, recent advancements in Deep Neural Networks (DNNs) open up new horizons for advanced analytics beyond simple statistical aggregation (Gardner and Dorling, 1998; Lian et al., 2018; Cheng et al., 2020; Cai et al., 2021).
@@
At its core, exploiting DNNs for advanced structured data analytics comprises two main phases: training and inference (Gardner and Dorling, 1998). The former primarily involves the construction of a DNN model and the training of this model on targeted data, while the latter utilizes the trained model to make predictions on new data. Notably, to deliver advanced DNN-driven analytics for informed decision-making, effectiveness, and efficiency are the two most important metrics to optimize for (Davis and Goadrich, 2006; Kwon et al., 2019; Cai et al., 2019). Specifically, effectiveness focuses on the inference phase, measuring the extent to which the predictions delivered by the model are accurate. Meanwhile, efficiency evaluates the requirements of the model in terms of response time and computational resources in both phases (Kwon et al., 2019).
@@
In real-world scenarios, analysts are often more interested in performing analytics on specific subsets of data. For instance, they may assess trends among patients diagnosed with a particular disease, or study behaviors of consumers in a certain age group. Consider the scenario illustrated in Figure 1, where an analyst examines the income situation within a specific demographic. This analysis informs critical decision-making, such as setting personal loan limits and adjusting local commodity prices. Naturally, the analyst seeks to build an effective predictive model, delivering accurate predictions for these subsets of tuples, meanwhile executing predictions efficiently with minimal response time and computational resources. However, there are two main challenges in achieving this objective.
@@
First, achieving efficient training for effective predictive modeling across analyst-specified subdatasets is challenging. Conventionally, a single general model is trained to support inference across all data tuples (He and Chua, 2017; Khamis et al., 2020; Gardner and Dorling, 1998). This approach is efficient and requires training only one model. However, such a model, optimized to capture the common patterns and general behaviors of the whole dataset, is likely not as effective in providing accurate predictions as a dedicated model trained on a specific subdataset of interest. Taking the scenario in Figure 1 as an example, a model trained explicitly for the group of men living in NYC would probably identify finer-grained patterns and behaviors pertinent to this subdataset. Given sufficient training instances, this dedicated model could outperform the general model significantly. Nonetheless, training a separate model for each subdataset is computationally prohibitive due to the combinatorial nature of potential subdatasets.
Secondly, enhancing the efficiency of inference execution for subdatasets queried from the RDBMS is also challenging from a system perspective. A major challenge is the coordination of the RDBMS data processing and the execution of inference tasks. Many existing solutions support the inference process using two separate systems (Wang et al., 2018; Crankshaw et al., 2017; Zhang et al., 2021), which requires transferring the inference data from an RDBMS to the inference system. However, this process is time-consuming, susceptible to errors, and may violate privacy and security requirements (Xu et al., 2022). More critically, additional data transfer overhead is introduced to adversely affect inference efficiency.
@@
To address the above challenges, we propose a novel SQL-awarE dynAmic moDel Slicing (LEADS) technique, which makes use of the meta-information in SQL queries to dynamically customize a predictive model, deriving meaningful insights in analysis tasks. Specifically, LEADS first constructs a high-capacity general model consisting of multiple replicas of the base model. These replicas, termed as experts in MoE, are trained to specialize in different problem subspaces for effective predictive modeling. To enhance effectiveness via MoE without incurring reduced inference efficiency, LEADS incorporates a SQL-aware gating network, which generates sparse gating weights based on the SQL query embedding to select necessary experts. Such a sliced model is trained to tailor for SQL queries and dedicated to the specified subdataset for inference effectiveness while maintaining efficiency. To further improve efficiency, we build an extension to integrate the inference process within the database via User-Defined-Function (UDF) with three optimizations: efficient execution allocation, memory sharing, and state caching. These approaches obviate the data transfer in separate systems, reduce data copying overhead, and reduce the cost associated with frequent model loading. We summarize our main contributions as follows.
@@
  • We formulate the SQL-aware structured data analytics problem, which requires efficient and effective predictive modeling on subdatasets specified by SQL queries. To the best of our knowledge, this is the first work for solving the problem.
@@
  • We propose a novel SQL-aware dynamic model slicing technique LEADS, which scales up the modeling capacity by replicating the base model as experts and devises a SQL-aware gating network to dynamically customize models for analytics tasks on SQL-specified subdatasets.
  • We implement an in-database inference extension on PostgreSQL to support SQL-aware structured data analytics, which incorporates three optimizations to improve inference efficiency.
@@
  • We conduct extensive experiments on five real-world datasets, which confirms the effectiveness of LEADS, with up to $3.95\%$ improvement in AUC for given data workloads, while the inference extension achieves up to $2.06$x speedup in terms of efficiency compared with the baseline approaches.
We have integrated LEADS with the extension onto NeurDB (Ooi et al., 2024; Zhao et al., 2025), our ongoing AI-powered autonomous data system implementation.
Figure 1. A workflow of supporting in-database analytics on income using SQL-aware dynamic model slicing.
@@
The paper is structured as follows. We introduce preliminaries in Section 2. We present details of LEADS in Section 3 and discuss the implementation in Section 4. Experimental results are presented in Section 5. We review related work in Section 6 and conclude in Section 7.
@@ 2. Preliminaries
In this section, we introduce the preliminaries of predictive modeling on structured data and present two key techniques central to our system, namely Mixture of Experts (MoE) for scaling up the model capacity while maintaining its inference efficiency via conditional computation (Shazeer et al., 2017), and sparsemax (Martins and Astudillo, 2016) for the informed selection of active experts to enhance efficiency. We also formally define the research problem, referred to as SQL-aware predictive modeling. Scalars, vectors, and matrices in the following part are denoted by $x$, $\mathbf{x}$ and $\mathbf{X}$ respectively.
Structured Data (or relational data, tabular data)  (Arasu and Garcia-Molina, 2003; BakIr, 2007) refers to the type of data that can be represented as tables. It is typically stored in a series of tables (relations) $\{\mathbf{T_{1}},\mathbf{T_{2}},...\}$ of columns and rows, which can be retrieved from a relational database with SQL operations, e.g., the projection, join, and aggregation. Tables in structured data are interconnected via foreign key attributes, where the value of a column in one table points to a unique row in another table. Technically, the structured data can be viewed as a logical table T comprising $N$ rows and $M$ columns. Each row is a tuple ${\mathbf{x}}=(x_{1},x_{2},...,x_{M})$, serving as a feature vector, where $x_{i}$ is the value of the $i$-th attribute and can be either numerical or categorical.
@@
Mixture of Experts (MoE)  (Jordan and Jacobs, 1994; Waterhouse et al., 1995; Shazeer et al., 2017) is a general ensemble learning and conditional computation technique to scale up the modeling capacity without incurring much computational overhead. MoE is particularly effective when the data exhibits complex patterns or variations (Riquelme et al., 2021; Fedus et al., 2022; Zhou et al., 2022). There are two main components in an MoE layer: expert models and a gating network. Expert models can be composed of homogeneous models that share the same architecture. They are adopted to divide problem space into different regions, where each expert specializes in handling a certain sub-region. The gating network outputs a set of gating weights for a given input, dynamically determining weights assigned to respective experts.
@@
Denoting the gating weights and outputs of experts as $\mathbf{w}=[w_{1},w_{2},...,w_{K}]$ and $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2}...,\mathbf{h}_{K}]$ respectively, where $K$ is the number of experts and $\mathbf{h}_{i}$ is the output of the $i$-th expert. the output of the MoE is a weighted average of these experts: $\hat{\mathbf{y}}=\sum_{i=1}^{K}w_{i}\mathbf{h}_{i}$. During training, the MoE model optimizes the gating network and experts simultaneously. The gating network learns to assign appropriate weight to experts, while the experts learn to make accurate predictions within their respective regions of expertise.
@@
MoE has found extensive application in various domains, notably in the large language model GPT-4 (OpenAI, 2023) for texts and the large vision-language model MoE-LLaVA (Lin et al., 2024) for images, which combines the benefits of large model capacity with efficient computation, by only engaging a fraction of the model parameters for each input. In LEADS, we apply the MoE technique to structured data, intending to harness its scalable modeling capacity for enhanced predictive effectiveness and efficiency.
@@
Sparsemax (Martins and Astudillo, 2016; Peters et al., 2019; Cai et al., 2021). Softmax transformation is a crucial function in the gating network, which maps an input vector $\mathbf{z}$ into a probability distribution $\mathbf{p}$ whose probabilities correspond proportionally to the exponential of its input values, i.e., $\mathrm{softmax(z_{j})}=\frac{\mathrm{exp(z_{j})}}{\mathrm{\sum_{i}exp(z_{i})}}$. The output can be denoted as the class probabilities or weights indicating the importance of corresponding inputs.
The softmax function is extensively used in DNNs due to its differentiable and convex properties. However, the output probabilities of the softmax function are dense, leading to less interpretability and effectiveness (Goodfellow et al., 2016; Cai et al., 2021). To overcome this limitation, $\alpha$-entmax is proposed to generalize the dense and sparse softmax, offering a parameter to adjust the sparsity level of probability distribution. Specifically, given a d-dimension probability as $\rm\Delta^{d}:=\{\mathbf{p}\in\mathbb{R}^{d}:\mathbf{p}\geq 0.\text{ }||\mathbf{p}||_{1}=1\}$, $\alpha$-entmax (Peters et al., 2019) can be interpreted in the variational form with Tsallis $\alpha$-entropies $\rm H_{\alpha}(\mathbf{p})$ (Tsallis, 1988):
@@
(1) $\rm\alpha\text{-}entmax(\mathbf{z})=\mathop{argmax}\limits_{\mathbf{p}\in\Delta^{d}}{\mathbf{p}^{T}\mathbf{z}+H_{\alpha}(\mathbf{p})}$
where $\mathrm{H_{\alpha}(\mathbf{p})}=-\frac{1}{\alpha(\alpha-1)}\sum_{j}(p_{j}-p^{\alpha}_{j})$ if $\rm\alpha\neq 1$, else $\rm H_{1}(\mathbf{p})=-\sum_{j}p_{j}\log p_{j}$, the Shannon entropy. A larger $\alpha$ causes $\alpha$-entmax to assign more zero in the probability distribution. Sparsemax is adopted in LEADS to sparsify the activated experts in MoE to enhance predictive efficiency.
@@
Problem Formulation. In structured data analytics, data analysts typically focus on specific subdataset characterized by shared attributes. For example, analysts may assess the readmission rates among patients diagnosed with a certain disease, or predict the e-commerce click-through rate (CTR) within a particular age group. Typically, for advanced analytical tasks that involve prediction, WHERE statement in a SQL query is executed first to select relevant tuples, to which DNNs are applied subsequently for inference. We refer to this process as SQL-aware predictive modeling.
@@
Given structured data $\mathbf{T}$ as illustrated in Section 2 and a SQL query denoted by $q$, there are two main steps in SQL-aware predictive modeling: data selection and model prediction. Corresponding to relational algebra, a generalized SQL query selection $q$ is expressed as $\sigma_{\varphi}(\textbf{T})$, where $\sigma$ is the unary operator for selection and $\varphi$ is the propositional formula in $q$. Typically, $\varphi$ consists of multiple predicates connected by logical operators. The selection $\sigma_{\varphi}(\textbf{T})$ retrieves all tuples in table T that satisfies $\varphi$, formally defined as $\sigma_{\varphi}(\textbf{T})=\{\textbf{x}:\textbf{x}\in\textbf{T},\varphi(\textbf{x})\}$. For simplicity, the subdataset retrieved by the SQL query $q$ is denoted as $\textbf{T}_{\varphi}=\{\textbf{x}_{1},\textbf{x}_{2},\cdots,\textbf{x}_{n}\}$, where $n$ is the number of tuples. Each tuple $\textbf{x}_{i}\in\mathbb{R}^{M}$ comprises $M$ attributes, and $\textbf{x}_{i}$ can be represented as a vector of features, i.e., $\textbf{x}_{i}=[x_{i,1},x_{i,2},\cdots,x_{i,M}]$. DNNs are then applied to perform prediction on these selected tuples, e.g., to predict the labels $\textbf{y}=\{y_{1},y_{2},\cdots,y_{n}\}$, aiming to derive meaningful insights, such as patients readmission rates in healthcare analytics or CTR in e-commerce. Technically, SQL-aware predictive modeling refers to making predictions on a selected subset of tuples retrieved from a logical table T based on a SQL query $q$ with a propositional formula $\varphi$.